OpenStudy (ray1998):

Hey guys! I really need help with this Algebra question if that's okay. I want you to walk me through it, because I am so lost. This is for Algebra 2. 1. Solve the system by Elimination. x + 5y - 4z = -10 2x - y + 5z = -9 2x - 10y - 5z = 0 A thank you in advance and whomever helps will receive medal :)

OpenStudy (ray1998):

@bibby

OpenStudy (bibby):

for #2, you can just substitute z=5 into the other 2 equations \(2x - y + z = -4 \implies 2x-y+5=-4\\2x-y=-9\) \(-2x + 3y - z = -10 \implies -2x + 3y - 5 = -10\\-2x + 3y = -5\) now our system looks like \(2x-y=-9\\-2x+3y=-5\) now we can solve one of the equations for a variable and the substitute to find the solution \(2x-y=-9\\2x=y-9\\2x+9=y\) replace that into the 2nd equation \(-2x+3y=-5\\-2x+3(2x+9)=-5\\-2x+6x+27=5\\4x=-32\\x=-8\) replace that into either equation to find y \(2x-y=-9 \\-8*2-y=-9 \\-16-y=-9 \\-y=7 \\y=-7 \)

OpenStudy (ray1998):

Thank you for helping on problem two :) @bibby

OpenStudy (bibby):

2x - y + 5z = -9 + 2x - 10y - 5z = 0 ---------------- 4x-11y=-9 having trouble doing the first one

OpenStudy (ray1998):

That's okay, thank you for trying, @bibby :)